-2m-8+12 find value of m.pls tell the answer along with explanation
 

Grade:6

1 Answers

Arun
6 years ago
Dear Sakshi
 
as
 – 2m – 8 + 12 = 0
Now
 – 2m + 4 = 0
now
 
2m = 4
m = 4/2
m = 2
 
Regards
Arun (askIITians forum expert)